# Comparative Performance of Collaborative Bandit Algorithms: Effect of Sparsity and Exploration Intensity

### v1

@@ 18       $\widehat{\Theta}_{t+1}\leftarrow\textup{mat}_{d,M}(\mathbf{A}_{t+1}^{-1}\mathbf{b}_{t+1})$
As clusters are intertwined in CoLin, that is, the collaboration structure allows users to influence other users, we can no longer let model parameters to be cluster-specific without introducing further unnecessary complexity. Hence, we time-index the parameters, e.g., $\mathbf{A}_{\mathrm{C}_{i}}$ (per cluster) becomes $\mathbf{A}_{t}$ (per time). This is enabled through introducing a block structure to permit cross relationships between items and users.
Before we introduce CoLin and FactorUCB, we define three special functions: $\textup{h}(\cdot)$ is the self-outer-product function, i.e., $\textup{h}(x)=x\otimes_{\textup{outer}}x$. Note that $\otimes$ is the standard Kronecker product for matrices which outputs a block matrix. The operation $\textup{mat}_{r,c}(\mathbf{x})$ on the vector $\mathbf{x}\in\mathbb{R}^{rc\times 1}$ outputs an $r\times c$ matrix. $\tilde{x}$ pads a row vector $x$ with preceding and/or trailing zeros as needed. [^10] In Algorithms 4 and 5, $\tilde{\bm{x}}_{a,u}$ is a row vector of dimension $1\times dM$, obtained by adding $d$-many preceding and trailing zeros for each user before and after user $u_{t}$, respectively ($(d-1)M$ zeros are added in total).
1 Input: $\mathcal{A}_{0}$, $\mathbf{W}\in\mathbb{R}^{M\times M}$, $\alpha_{1}=\alpha_{2}=0.375$.
@@ 20       $\mathbf{\hat{v}}_{a,t+1}\leftarrow\mathbf{E}_{a_{t},t+1}^{-1}\mathbf{d}_{a_{t},t+1}$
We present CoLin in Algorithm 4 and FactorUCB in Algorithm 5. In these two algorithms, we restrict the number of items available in each time period, i.e., $|\mathcal{A}_{t}|=K$ $\forall t\in[T]$, for some known and time-invariant $K$.
Comparing M-LinUCB in Algorithm 3 to CoLin in Algorithm 4, we observe the introduction of the so-called similarity matrix $\mathbf{W}$. This matrix relates the estimated bandit parameter of each user cluster to that of other user clusters. An element $(i,j)$ of the similarity matrix $\mathbf{W}$ governs how much the bandit parameter associated with user cluster $i$ affects the bandit parameter associated with user cluster $j$. That is, column $j$ of $\mathbf{W}$ contains all the information necessary to propagate the user preferences from all other clusters $i\neq j$. Therefore, the similarity matrix $\mathbf{W}$ is column-normalized, and its elements quantify how similar one user cluster is to other user clusters. Since all elements of $\mathbf{W}$ are between non-negative column-wise sum up to zero, a bandit parameter for a cluster is the convex combination of all clusters, including itself.
Observe that in Line 12 of Algorithm 3, the UCB of an article is calculated by only using the bandit parameter for the current user’s cluster. In Line 11 of Algorithm 4, UCB is now calculated by generating a shared bandit parameter for all user clusters the current user has related preferences with. Similarly, the standard deviation term in Line 11 of Algorithm 4 now depends on the similarity matrix $\mathbf{W}$ as well (compare to that of in Line 12 of Algorithm 3).
Then, going from CoLin to FactorUCB in Algorithm 5, we observe two new concepts: (1) increase in the dimension of the bandit parameters by the dimension of latent factors ($d_{l}$) introduced to the items, increasing the dimensionality to $d+d_{l}$, and (2) introduce a learning scheme for latent item factors that make up the context for the items, i.e., item contexts are made up of two parts: static and dynamic. Although we combine the unknown parameters in $\widehat{\Theta}$, for calculating user specific bandit parameters and arm specific latent factors, we let $\widehat{\Theta}^{\bm{x}}$ denote the user part and $\widehat{\Theta}^{\mathbf{\hat{v}}}$ denote the arm (latent factor) part, i.e., $\widehat{\Theta}=\textup{concat}\left(\widehat{\Theta}^{\bm{x}},\widehat{\Theta}^{\mathbf{\hat{v}}}\right)$. [^11] We use $\hat{q}$ to denote that $q$ is a learnable parameter. In the statement of the algorithm, we drop “concat” and simply put two vectors within closed parentheses to represent they are being concatenated.
@@ 4 Numerical Experiments
We compare and contrast M-LinUCB (Algorithm 3), CoLin (Algorithm 4) and FactorUCB (Algorithm 5) algorithms on the Yahoo! Today Module (2009) dataset (version R6A). [^12] Obtained from https://webscope.sandbox.yahoo.com/catalog.php?datatype=r&guccounter=1 We first describe details of the experiments, and then comment on the effect of varying sparsity levels and exploration intensity.
We associate bandit parameters with users because we aim to model the similarity across users’ preferences rather than items’ characteristics. Moreover, the choice of associating bandit parameters with users is often more convenient as the items’ characteristics are under the control of the decision-maker, while the users’ characteristics are exogenous to the platform and harder to account for. Note that, in general, there are many considerations that may affect this choice, e.g., number of items or users and their ratios.
@@ 4.1 Offline Policy Evaluation and Sparsity
    Receive a user and a pool of candidate arms the users can receive as recommendation. For the Yahoo! Today Module dataset, this pool consists of 20 articles that are shown in an instance.
@@
We utilize users features to generate cluster centroids, i.e., assign each user to their closest cluster, and associate the clusters themselves with the bandit parameters. We use the cluster centroids to define the collaboration across clusters. The dot product of the centroid of cluster $i$ with that of cluster $j$ is used to obtain the intensity of the collaboration. We then normalize these values to generate the collaboration intensity and call them weights, each of which is a number between 00 and $1$.
We also control the sparsity in this collaboration structure. By controlling sparsity, we mean only preserving the top $x\%$ weights across users. In this case, we say that this instance has Sparsity $x\%$. If we introduce sparsity, i.e., set $x$ to a positive number strictly less than $100$, we normalize the weights before using them in the algorithms.
@@ 4.2 Results on Yahoo! Today Module Dataset
In Figures 1 and 2, we report rolling averages of CTR values. The average CTR value is calculated every $2,000$ time periods, generating the $22,905$ data points used to create these figures. Using a rolling average of $500$ to average the CTR values, we generate Figures 1 and 2. (Time, the $x$-axis, goes from zero to $22,405$.) In Figures 3, 4, and 5, we report the cumulative average CTR values.
@@ 4.2.1 Comparing Experiment Results by Varying the Cluster Size and the Sparsity
Figures 1 and 2 are organized by $\alpha$: Figure 1 has $\alpha=0.5$ and Figure 2 has $\alpha=1$. Recall that we normalize the performance of all algorithms by a random policy, where the random algorithm recommends each candidate article with the same probability in each period. In both figures, the figures on the left has cluster size $160$, and the figures on the right has cluster size $80$. We denote a cluster size $160$ as C$(160)$ and a sparsity of $100\%$ as S$(100)$, to refrain from continuously repeating those words. (We leave them as they are in the figure captions.)
@@
We observe that FactorUCB performs best for C$(160)$ and S$(50)$, with C$(160)$ and S$(25)$ being a close second: See Figure 5(a). FactorUCB with C$(160)$ and S$(50)$ outperforms the random policy by $67.5\%$, and the best CoLin by $24.6\%$ (see Figure 3(a)). The best CoLin outperforms the best M-LinUCB by $7\%$. From this, we can immediately reason that the marginal value of the similarity matrix $\mathbf{W}$ is low since the improvement of CoLin over M-LinUCB is fairly small compared to improvement of FactorUCB over M-LinUCB. Especially, CoLin with C$(80)$ and S$(100)$ frequently under-performs M-LinUCB in Figure 2(b) (and its cumulative CTR average eventually dips below that of M-LinUCB, see Figure 3(b)).
@@
CoLin improves for C$(80)$ when the sparsity is reduced, or the cluster size is increased with any sparsity level. This suggests that a cluster size of $80$ with full sparsity is of low quality. Hence, we do not test a cluster size smaller than $80$ in our experiments. The original FactorUCB work (Wang et al., , 2017) also notes that, for the same Yahoo! Today Module dataset, FactorUCB with a trivial similarity matrix $\mathbf{W}$ (i.e., $\mathbf{W}=\mathbf{I}$) outperforms FactorUCB with a non-trivial $\mathbf{W}$. While we do not test that particular configuration, comparing CoLin to M-LinUCB helps in arriving to a similar observation. Therefore, introducing a static $\mathbf{W}$ only may not be worth for the performance improvement as it brings with a high computation complexity. We present another empirical proof for the potential misspecification of $\mathbf{W}$ in the next subsection.
@@
FactorUCB performs best for C$(160)$ when S$(50)$, and for C$(80)$ it performs the best when S$(100)$. This suggests that, while a larger cluster size with full connectivity may degrade the performance, a smaller cluster size with any sparsity level is not a match for a larger cluster size. (See Figure 5(a) for a clear picture.) Therefore, we argue that introducing sparsity and allowing a loosely connected forest (or clique) structure with large cluster sizes generate the best improvement, bringing about an improvement of $2.5\%$ over its smaller cluster size counterpart. This is consistent with intuition: A larger cluster size means more flexibility in modeling the variety, but full connectedness may force propagation of information irrelevant to some, taking away from those who would benefit more from information propagation. [^14] A larger cluster size with smaller sparsity, heuristically, only preserves the strong-enough relationships. Coincidentally, we observe that the performance improves by $2.3\%$ when sparsity is halved for C$(160)$ from S$(100)$ to S$(50)$.
@@ 4.2.2 Comparing Experiment Results by Varying $\alpha$
In Figure 1, CoLin performs best for C$(80)$ and S$(50)$: It outperforms the random policy by $34.4\%$. When $\alpha$ is increased from $0.5$ to $1$ in Figure 2, CoLin performs best for C$(160)$ and S$(50)$: It outperforms the random policy by $33.9\%$. This difference can be seen clearly in Figure 5(d). On average, CoLin outperforms the random policy by $30.6\%$ and $31.2\%$ when $\alpha=0.5$ and $\alpha=1$, respectively. That is, while the best performance is when $\alpha=0.5$, so is the worst performance. When $\alpha=1$, the performance is more stable across different cluster sizes and sparsity values. This may suggest that the similarity matrix $\mathbf{W}$ is slightly misspecified and tendency to explore more is acting as a correction, reducing the variance due to hyperparameter selection.
@@ 4.2.3 Varying the Exploration Intensity in FactorUCB
We close the discussion on numerical results in this section by presenting the results of additional tests conducted on the selection of $\alpha_{1}$ and $\alpha_{2}$ parameters for FactorUCB. We run two additional tests where we compare the original setting of $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ with the following two settings: (1) $\alpha_{1}=2\alpha_{2}=0.5$, and, (2) $0.52\alpha_{1}=\alpha_{2}=$ 0.5. We set cluster size to $160$ and sparsity to $50\%$.
@@
The original instance at $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ with cluster size $160$ and sparsity $50\%$ performs the best among the earlier FactorUCB results, outperforming the random policy by $67.5\%$. The instance with $\alpha_{1}=2\alpha_{2}=0.5$ outperforms the random policy by $65.2\%$, and the instance with $0.52\alpha_{1}=\alpha_{2}=$ 0.5 outperforms the random policy by $62.2\%$.
In Figure 6(a), we can see that the $0.52\alpha_{1}=\alpha_{2}=$ 0.5 instance lags behind most of the time while the $\alpha_{1}=2\alpha_{2}=0.5$ instance matches the original instance’s behavior more closely. Although, in Figure 6(b), we observe that the $\alpha_{1}=2\alpha_{2}=0.5$ instance, while getting very close to the original instance around time $15,000$, eventually diverges. These results confirm our setting of $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ for FactorUCB. For the Yahoo dataset, we achieve the best performance out of FactorUCB using equal $\alpha$ values.
@@ 5 Conclusion
Although a temporal design can remedy the fast-population-shift cases, we already observe FactorUCB is high variance, as discussed in Section 4.2.4. We conjecture that this is due to two structural elements of the FactorUCB algorithm: (1) latent factors are learned through interactions with historical data, and hence, are high variance; and (2) latent factors are not as helpful for brand new articles with insufficient information. For the first point, we conjecture that controlling the change (i.e., smoothing) in latent parameter updates can help. And, for the second point, designing a distinct similarity matrix $\mathbf{W}$ for the articles may be helpful as article features may be used to extract similarity values.181818 Similarly, in Section 4.2.3, we discuss that varying the $\alpha$ values can be a way to control the potential misspecification in $\mathbf{W}$, and identifying the rules for making parameter updates to that end is an exciting future research direction.
@@ Acknowledgement
We would like to thank Sharon O’Shea Greenbach (Sr. Counsel & Director, Regulatory Policy at DFS) for carefully reviewing the manuscript and their helpful comments. The author expresses gratitude for the mentorship of Ashkan Golgoon and Arjun Ravi Kannan during his internship at Discover Financial Services, which greatly contributed to the success of this work. The views and opinions expressed in this paper are solely those of the author and do not necessarily reflect the official policy or position of their respective employers or any affiliated organizations. Any opinions, findings, conclusions, or recommendations are those of the author alone.

### v2

@@ 18  $\widehat{\Theta}_{t+1}\leftarrow\textup{mat}_{d,M}(\mathbf{A}_{t+1}^{-1}\mathbf{b}_{t+1})$
As clusters are intertwined in CoLin, that is, the collaboration structure allows users to influence other users, we can no longer let model parameters be cluster-specific without introducing further unnecessary complexity. Hence, we time-index the parameters, e.g., $\mathbf{A}_{\mathrm{C}_{i}}$ (per cluster) becomes $\mathbf{A}_{t}$ (per time). This is enabled through introducing a block structure to permit cross relationships between items and users.
Before we introduce CoLin and FactorUCB, we define three special functions: $\textup{h}(\cdot)$ is the self-outer-product function, i.e., $\textup{h}(x)=x\otimes_{\textup{outer}}x$. Note that $\otimes$ is the standard Kronecker product for matrices which outputs a block matrix. The operation $\textup{mat}_{r,c}(\mathbf{x})$ on the vector $\mathbf{x}\in\mathbb{R}^{rc\times 1}$ outputs an $r\times c$ matrix. $\tilde{x}$ pads a row vector $x$ with preceding and/or trailing zeros as needed. [^10] In Algorithms 4 and 5, $\tilde{\bm{x}}_{a,u}$ is a row vector of dimension $1\times dM$, obtained by adding $d$-many preceding and trailing zeros for each user before and after user $u_{t}$, respectively ((d1)M(d-1)M zeros are added in total).
1Input: $\mathcal{A}_{0}$, $\mathbf{W}\in\mathbb{R}^{M\times M}$, $\alpha_{1}=\alpha_{2}=0.375$.
@@ 20  $\mathbf{\hat{v}}_{a,t+1}\leftarrow\mathbf{E}_{a_{t},t+1}^{-1}\mathbf{d}_{a_{t},t+1}$
We present CoLin in Algorithm 4 and FactorUCB in Algorithm 5. In both algorithms, we restrict the number of items available in each time period, i.e., $|\mathcal{A}_{t}|=K$ $\forall t\in[T]$, for some known and time-invariant $K$.
Comparing M-LinUCB in Algorithm 3 to CoLin in Algorithm 4, we observe the introduction of the so-called similarity matrix $\mathbf{W}$. This matrix relates the estimated bandit parameter of each user cluster to that of other user clusters. An element $(i,j)$ of the similarity matrix $\mathbf{W}$ governs how much the bandit parameter associated with user cluster $i$ affects the bandit parameter associated with user cluster $j$. That is, column $j$ of $\mathbf{W}$ contains all the information necessary to propagate the user preferences from all other clusters $i\neq j$. The similarity matrix $\mathbf{W}$ is column-normalized and non-negative. This implies that a bandit parameter for a cluster is defined as the convex combination of all clusters’ bandit parameters, including its own.
Observe that in Line 12 of Algorithm 3, the UCB of an article is calculated by only using the bandit parameter for the current user’s cluster. In Line 11 of Algorithm 4, UCB is now calculated by generating a shared bandit parameter for all user clusters the current user has related preferences with. Similarly, the standard deviation term in Line 11 of Algorithm 4 now depends on the similarity matrix $\mathbf{W}$ as well (compare to that in Line 12 of Algorithm 3).
[…]
@@ 4 Numerical Experiments
We compare and contrast M-LinUCB (Algorithm 3), CoLin (Algorithm 4) and FactorUCB (Algorithm 5) algorithms on the Yahoo! Today Module (2009) dataset (version R6A). [^12] Obtained from https://webscope.sandbox.yahoo.com/catalog.php?datatype=r&guccounter=1 First, we detail our experiments, then we discuss how varying sparsity and exploration intensity affect the results.
We associate bandit parameters with users because we aim to model the similarity across users’ preferences rather than items’ characteristics. Moreover, the choice of associating bandit parameters with users is often more convenient as the items’ characteristics are under the control of the decision-maker, while the users’ characteristics are exogenous to the platform and harder to account for. Note that, in general, many considerations may affect this choice, e.g., number of items or users and their ratios.
@@ 4.1 Offline Policy Evaluation and Sparsity
    Receive a user and a pool of candidate arms the user can receive as recommendation. For the Yahoo! Today Module dataset, this pool consists of 20 articles that are shown in an instance.
@@
We utilize user features to create cluster centroids, assigning each user to the nearest cluster. We then associated these clusters with the bandit parameters. We use the cluster centroids to define the collaboration across clusters. The dot product of the centroid of cluster $i$ with that of cluster $j$ is used to obtain the intensity of the collaboration. We then normalize these values to generate the collaboration intensity and call them weights, each of which is a number between 0 and $1$.
We also control the sparsity in this collaboration structure. By controlling sparsity, we mean only preserving the top $x\%$ weights across users. In this case, we say that this instance has Sparsity $x\%$. If we introduce sparsity (i.e., set $x$ to a positive number strictly less than $100$), we normalize the weights before using them in the algorithms.
@@ 4.2 Results on Yahoo! Today Module Dataset
In Figures 1 and 2, we report rolling averages of CTR values. We calculated the average CTR every 2,000 time periods. This process generated 22,905 data points, which we used to create the figures. Using a rolling average of $500$ to average the CTR values, we generate Figures 1 and 2. (Time, the $x$-axis, goes from zero to $22,405$.) In Figures 3, 4, and 5, we report the cumulative average CTR values.
@@ 4.2.1 Comparing Experiment Results by Varying the Cluster Size and the Sparsity
Figures 1 and 2 are organized by $\alpha$: Figure 1 has $\alpha=0.5$ and Figure 2 has $\alpha=1$. Recall that we normalize the performance of all algorithms by a random policy, where the random algorithm recommends each candidate article with the same probability in each period. In both figures, the figures on the left has cluster size $160$, and the figures on the right has cluster size $80$. We denote a cluster size $160$ as C(160)(160) and a sparsity of $100\%$ as S(100)(100), to refrain from continuously repeating those words. (We leave them as they are in the figure captions.)
@@
We observe that FactorUCB performs best for C(160)(160) and S(50)(50), with C(160)(160) and S(25)(25) being a close second: See Figure 5(a). FactorUCB with C(160)(160) and S(50)(50) outperforms the random policy by $67.5\%$, and the best CoLin by $24.6\%$ (see Figure 3(a)). The best CoLin outperforms the best M-LinUCB by $7\%$. From this, we can immediately reason that the marginal value of the similarity matrix $\mathbf{W}$ is low since the improvement of CoLin over M-LinUCB is fairly small compared to improvement of FactorUCB over M-LinUCB. Especially, CoLin with C(80)(80) and S(100)(100) frequently under-performs M-LinUCB in Figure 2(b) (and its cumulative CTR average eventually dips below that of M-LinUCB, see Figure 3(b)).
@@
CoLin improves for C(80)(80) when the sparsity is reduced, or the cluster size is increased with any sparsity level. This suggests that a cluster size of $80$ with full sparsity is of low quality. Hence, we do not test a cluster size smaller than $80$ in our experiments. The original FactorUCB work (Wang et al., , 2017) also notes that, for the same Yahoo! Today Module dataset, FactorUCB with a trivial similarity matrix $\mathbf{W}$ (i.e., $\mathbf{W}=\mathbf{I}$) outperforms FactorUCB with a non-trivial $\mathbf{W}$. While we do not test that particular configuration, comparing CoLin to M-LinUCB helps in arriving to a similar observation. Therefore, introducing a static $\mathbf{W}$ only may not be worth for the performance improvement as it brings with a high computation complexity. We present another empirical proof for the potential misspecification of $\mathbf{W}$ in the next subsection.
@@
FactorUCB performs best for C(160)(160) when S(50)(50), and for C(80)(80) it performs the best when S(100)(100). This suggests that, while a larger cluster size with full connectivity may degrade the performance, a smaller cluster size with any sparsity level is not a match for a larger cluster size. (See Figure 5(a) for a clear picture.) Therefore, we argue that introducing sparsity and allowing a loosely connected forest (or clique) structure with large cluster sizes generate the best improvement, bringing about an improvement of $2.5\%$ over its smaller cluster size counterpart. This aligns with our intuition: while a larger cluster size allows for more flexible modeling, full connectivity might spread irrelevant information to some users. This could hurt the performance for users who would benefit from more targeted information sharing. [^14] A larger cluster size with smaller sparsity, heuristically, only preserves the strong-enough relationships. Coincidentally, we observe that the performance improves by $2.3\%$ when sparsity is halved for C(160)(160) from S(100)(100) to S(50)(50).
@@ 4.2.2 Comparing Experiment Results by Varying $\alpha$
In Figure 1, CoLin performs best for C(80)(80) and S(50)(50): It outperforms the random policy by $34.4\%$. When $\alpha$ is increased from $0.5$ to $1$ in Figure 2, CoLin performs best for C(160)(160) and S(50)(50): It outperforms the random policy by $33.9\%$. This difference can be seen clearly in Figure 5(d). On average, CoLin outperforms the random policy by $30.6\%$ and $31.2\%$ when $\alpha=0.5$ and $\alpha=1$, respectively. That is, while the best performance is when $\alpha=0.5$, so is the worst performance. When $\alpha=1$, the performance is more stable across different cluster sizes and sparsity values. This may suggest that the similarity matrix $\mathbf{W}$ is slightly misspecified and tendency to explore more is acting as a correction, reducing the variance due to hyperparameter selection.
@@ 4.2.3 Varying the Exploration Intensity in FactorUCB
We close the discussion on numerical results in this section by presenting the results of additional tests conducted on the selection of $\alpha_{1}$ and $\alpha_{2}$ parameters for FactorUCB. We run two additional tests where we compare the original setting of $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ with the following two settings: (1) $\alpha_{1}=2\alpha_{2}=0.5$, and, (2) $2\alpha_{1}=\alpha_{2}=0.5$. We set cluster size to $160$ and sparsity to $50\%$.
@@
The original instance at $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ with cluster size $160$ and sparsity $50\%$ performs the best among the earlier FactorUCB results, outperforming the random policy by $67.5\%$. The instance with $\alpha_{1}=2\alpha_{2}=0.5$ outperforms the random policy by $65.2\%$, and the instance with $2\alpha_{1}=\alpha_{2}=0.5$ outperforms the random policy by $62.2\%$.
In Figure 6(a), we can see that the $2\alpha_{1}=\alpha_{2}=0.5$ instance lags behind most of the time while the $\alpha_{1}=2\alpha_{2}=0.5$ instance matches the original instance’s behavior more closely. Although, in Figure 6(b), we observe that the $\alpha_{1}=2\alpha_{2}=0.5$ instance, while getting very close to the original instance around time $15,000$, eventually diverges. These results confirm our setting of $\alpha_{1}$ = $\alpha_{2}$ = $0.375$ for FactorUCB. For the Yahoo dataset, we achieve the best performance out of FactorUCB using equal $\alpha$ values.
@@ 5 Conclusion
Although a temporal design can remedy the fast-population-shift cases, we already observe FactorUCB is high variance, as discussed in Section 4.2.4. We conjecture that this is due to two structural elements of the FactorUCB algorithm: (1) latent factors are learned through interactions with historical data, and hence, are high variance; and (2) latent factors are not as helpful for brand new articles with insufficient information. For the first point, we conjecture that controlling the change (i.e., smoothing) in latent parameter updates can help. And, for the second point, designing a distinct similarity matrix $\mathbf{W}$ for the articles may be helpful as article features may be used to extract similarity values. [^18] Similarly, in Section 4.2.3, we discuss that varying the $\alpha$ values can be a way to control the potential misspecification in $\mathbf{W}$, and identifying the rules for making parameter updates to that end is an exciting future research direction.
@@ Acknowledgement
The first author expresses gratitude for the mentorship of Arjun Ravi Kannan during his internship at Discover Financial Services, which greatly contributed to the success of this work. The views and opinions expressed in this paper are solely those of the authors and do not necessarily reflect the official policy or position of their respective employers or any affiliated organizations. Any opinions, findings, conclusions, or recommendations are those of the authors alone.